OpenStudy (erniellyd28):

a rectangular garden is 47 feet longer than it is wide. its 278 feet (a. let w be the width of the garden and l be the length. write a system of linear equations whose solution gives the width and length of the garden b. use the system of substitution to solve the system.)

OpenStudy (zzr0ck3r):

its 278 feet?

OpenStudy (erniellyd28):

thats as far as i got

OpenStudy (zzr0ck3r):

? You wrote "its 278 feet" What is 278 feet?

OpenStudy (erniellyd28):

2l+2w=278,l-w=47

OpenStudy (zzr0ck3r):

its perimeter is 278 feet?

OpenStudy (erniellyd28):

yes

OpenStudy (zzr0ck3r):

ok, you are almost there. Can you solve the second equation for l?

OpenStudy (erniellyd28):

thats where i get lost

OpenStudy (zzr0ck3r):

\(l-w=47\) add \(w\) to both sides and what do you get?

OpenStudy (erniellyd28):

im very sorry but i have no clue